# Transformer Alignment in Large Language Models

### v1

@@ 4.1 Alignment of Singular Vectors of Residual Jacobians
To understand this phenomenon more precisely, for each LLM, we measure the average alignment (which we define to be negative misalignment) across prompts in the 6 evaluation datasets (Section 3.2), where for each prompt, the misalignment $m(A_{i,j,K})$ is averaged over layers $i,j=1,\ldots,L$. We plot the negative misalignment values against the benchmark scores across several LLMs (Figure 9). Our results reveal a consistent pattern: an increase in alignment corresponds to an increase in benchmark scores. This observation suggests a compelling relationship between stronger alignment of singular vectors of Residual Jacobians and improved generalization performance.
Figure 9: Relationship between Alignment and Benchmark Score. The plot illustrates the relationship between average alignment (negative misalignment) and benchmark score across various LLMs. The left subplot includes models fine-tuned for coding and instruction, while the right excludes these models. We make this distinction as these fine-tuned models are for code and instruction, rather than being fine-tuned on the benchmark tasks. The data shoes a noticeable trend: increased alignment correlates with elevated benchmark scores. In particular, when excluding the code and instruction models, the linear regression gives an $R^{2}$ value of $0.7214$. This suggests that stronger singular vector alignment may correspond with enhanced generalization performance.
4.2 Linearity of Trajectories
@@
We observe (TA1) in each LLM, evident in the LSS of layer-wise trajectories at initialization and after training evaluated over benchmark prompts (Figure 10), in agreement with (RA1) in ResNets. Linear and expansive dynamical behavior is demonstrated in Llama-3 70B, Mistral 30B, and NeoX 20B through low dimensional projections of embedding trajectories (Figure LABEL:figs:pca). The LSS of the trajectories has an average value $4.25$ across trained models, while taking average values of $6.54$ at initialization, and is evidenced by the low variation across benchmark prompts (Figure 10).
@@
Figure 10: (TA1): Linearity of layer-wise trajectories. The figure depicts the line-shape score (LSS) of embedding trajectories, as discussed in Section 4.2), computed on 1,200 prompts of the HuggingFace Open LLM Leaderboard (Section 3.2) for a variety of trained (black) and randomly initialized (red) LLMs (Appendix A.1). Plotted are the median values over all prompts, and are accompanied with uncertainty intervals depicting the inter-quartile range of the results for each model. Models are sorted by increasing benchmark performance. Across every model, training increases linearity (lower LSS) of embedding trajectories when compared to that of untrained models (greater LSS).
After training, LSS appears largely independent of the parameter budget, embedding dimension, and the HuggingFace OpenLLM benchmark score. At initialization, the non-linearity of the trajectories grows with model size, embedding dimension, and number of layers (Figure 10). Moreover, the discrepancy between trained and untrained models increases with the number of parameters, hidden layers, and embedding dimension.
4.3 Layer-wise Growth
@@
Across all LLMs considered, the hidden trajectories exhibit distinctly exponential growth that emerges with training (Figure 11), displaying the second property of (TA1). This property, which arises as a result of Jacobian alignment, is distinct from (RA1) in the sense that distances between consecutive layers grow exponentially instead of linearly with depth, and highlights a key difference between the trajectories in ResNet classifiers and transformers. Most of the LLMs considered exhibit a low coefficient of variation across prompts, showing the robustness of this property across a variety of tasks. The expodistance measurements show little dependence on model parameters, and take on consistent and small values across the models (Figure 11).
@@
Figure 11: (TA1): Coefficient of variation of layer-wise exponent. The figure presents the coefficient of variation of the layer-wise exponent (Section 3.4) computed on 1,200 prompts of the HuggingFace Open LLM Leaderboard (Section 3.2) for trained LLMs (Appendix A.1). Plotted are the median values over all prompts, and are accompanied with uncertainty intervals depicting the inter-quartile range of the results for each model. LLMs are sorted by increasing benchmark performance, while the vertical axis has a base 10 logarithmic scale. The coefficient of variation for most models is low, suggesting that the growth of the embedding norms follows an exponential pattern across different layers, characterized by a relatively constant exponent.
At initialization, each model consistently displays equally (rather than exponentially) distanced trajectories, as reflected by the extremely low coefficients of variation in the norms of their layer-wise differences (Figure 23). Conversely, the trajectories of the trained models appear to be significantly less equidistant across layers, exhibiting substantially larger coefficients of variation with far less consistency across prompts. This behaviour, however, is expected and is discussed in more detail in Section 5.4.
@@ 5.1 Correlation with Generalization
As discussed in Section 4.1, stronger alignment of Residual Jacobians singular vectors is related to improved generalization (Figure 9). This relationship between (TA2) with LLM benchmark scores may assist in evaluating the training of transformers. Understanding the relationship between training methods and regularity in representations holds potential implications for the enhancement of LLMs. This finding suggests that development of techniques to amplify Transformer Alignment may lead to favourable model performance. The connection between hidden representation regularity, generalization, and LLM training, as prompted by our work, presents an intriguing research direction with important practical implications.
5.2 Regularity and Training
Many LLMs are based on the same transformer architecture, yet when trained by independent individuals and organizations, training directly affects the emergence of residual Jacobian alignment (Figure 9), presence of linearity (Figure 10), and extent of exponential growth (Figure 11). The effect of training is further evident when comparing base LLMs to their fine-tuned counterparts, since differences in measurements are a direct product of further optimization of the same weights. Our findings further highlight the important effect of training on the dynamical properties of hidden trajectories in transformers.
5.3 Comparison to ResNets
Linearity and Equidistance.
Linearity in hidden trajectories, as observed in ResNets (RA1) Li and Papyan [2023], Gai and Zhang [2021], also emerges with training in LLMs (TA1). The mean LSS value among the evaluated LLMs is $4.24$ (Figure 10), greater than LSS measurements observed for ResNets (Gai and Zhang [2021], page 18) which range between $2.0$-$3.0$. In both architectures, training induces improved linearity and regularity (Figure LABEL:figs:pca) in trajectories. In contrast to ResNets, trajectories are not equidistant, instead showing exponential growth between layers (Figure 11). A low coefficient of variation displays the presence of (TA1) on token trajectories, in analogy with the equidistant spacing of (RA1). In both ResNet classifiers and LLM transformers, there is an evident level of regularity in hidden representations (Figure LABEL:figs:pca).
Residual Jacobian Singular Vector Alignment.
The alignment of singular vectors of Residual Jacobians is present across LLMs (Figure 9), similar to how it occurs in ResNets [Li and Papyan, 2023], and does not occur at initialization.
Rank of Residual Jacobians.
Li and Papyan [2023] show that residual Jacobians have rank at most $C$, the number of classes. This analogous result automatically holds for LLMs since the vocabulary size is significantly greater than the embedding dimension of the transformer blocks.
Singular Value Scaling.
Li and Papyan [2023] observe that top singular values of residual Jacobians scale inversely with depth. In trained LLMs, however, top singular values do not show a consistent depth scaling across models (Figure LABEL:fig:svalues_trained), notably differing from classification ResNets. In addition, the distribution of singular values at each layer varies significantly between models. Interestingly, singular value scaling is actually more present in untrained transformers (FigureLABEL:fig:svalues_untrained).
5.4 Emergence of Exponential Growth
Trained ResNets distinctly feature equidistance in hidden trajectories, while measurements in transformers display a layer-wise exponential growth in norm. [Li and Papyan, 2023] prove that equidistance in certain classifiers emerges under the assumptions of (RA 2-4), in particular, the authors assume that top singular values of residual Jacobians scale inversely with model depth. In transformers, top singular values do not show a similar scaling, possibly suggesting that (RA4) is a necessary condition for equidistance.
Emergence of (TA1) appears to be a result of (TA2) and a relationship between the singular values, lying analogously to the emergence of (RA1) as a result of (RA 2-4). (TA2) implies that linearizations of hidden trajectories take a simple form after training, satisfying a linear difference equation (see Appendix A.2). In contrast to (RA4), the top singular values of the Jacobian blocks exhibit low variation across depth in some models (Figure LABEL:fig:svalues_trained). Under these assumptions, decomposing the matrix $A$ in Equation (16) as $A=U\Sigma V^{T}$, we obtain
$|x(t)|\approx\exp(\sigma t)|x(0)|,$ (14)
@@ 5.4 Emergence of Exponential Growth
where $\sigma$ is the top singular value. Since $\sigma$ is independent of depth, the above norm will be exponential in $t$, and directly explains the pattern seen in Figure 11.
Common in many models at initialization, the top singular values of the Jacobian matrices decay reciprocally with depth (Figure LABEL:fig:svalues_untrained). As a result, relatively linear trajectories appearing in the untrained models is not surprising. Randomly initialized Jacobian matrices with decaying singular values, coupled with Equation 17 results in the approximate exponential growth becoming linearized due to the receding impact of each layer.
@@ 6 Related Work
Residual Networks. ResNets have been viewed as an ensemble of shallow networks [Veit et al., 2016], with studies delving into the scaling behaviour of their trained weights [Cohen et al., 2021]. The linearization of residual blocks by their Residual Jacobians was first explored by Rothauge et al. [2019], who examined Residual Jacobians and their spectra in the context of stability analysis, and later by Li and Papyan [2023] who discovered Residual Alignment. We continue this line of work by investigating Residual Jacobians in transformer architectures and uncover a similar phenomenon in pretrained LLMs.
In-context learning. Previous works have observed that LLMs are able to perform in context learning [von Oswald et al., 2023, Bai et al., 2024, Ahn et al., 2023, Akyürek et al., 2023, Xie et al., 2021, Hahn and Goyal, 2023, Xing et al., 2024]: the model is capable of performing a particular machine learning task through a few examples provided in a single prompt or context. Subsequent theoretical works explained this phenomenon by postulating that intermediate representations of deep networks are capable of emulating classical learning algorithms like solving least squares through gradient descent, solving LASSO, and even training a shallow network.
Our work provides empirical evidence for these works by demonstrating that the transformer blocks of LLMs are coupled in a way that would only emerge if consecutive blocks jointly try to solve a specific problem.
@@
Neural ODE. Introduced by [Chen et al., 2018], Neural ODE view ResNets as a discretized dynamical process, with past work [Sander et al., 2022] showing the convergence of Residual Networks to a system of linear ODE. [Zhong et al., 2022, Li et al., 2021] studied neural ODE specific perspective to transformers and other LLM architectures. In our work, the presence of residual connections and emergence of Jacobian Alignment in transformers suggests that a discretization of a simple ODE emerges in LLMs.
@@
Information Processing in LLMs. Previously, Katz and Belinkov [2023], Bietti et al. [2023] have described memory and semantic flow in intermediate states of transformers while van Aken et al. [2019] have examined BERT hidden trajectories. Investigation of LLMs for reasoning tasks and interpretability remains an important research focus [Huang and Chang, 2023].
@@
Hidden Representation Dynamics. Previous studies [Gai and Zhang, 2021, Haber and Ruthotto, 2017, Ee, 2017], have focused on interpreting deep neural networks from a dynamical systems perspective. In particular, the layer-wise transformations between transformer blocks are interpreted as discrete approximations to continuous curves through representation space. Throughout training, numerical experiments have shown that both plain networks and ResNets are aptly approximated by a geodesic curve in Wasserstein space [Gai and Zhang, 2021]. Other interpretations involve partitioning activation space into basins of attraction, with fixed points representing discrete thoughts [Nam et al., 2023]. For further related works, see [Geshkovski et al., 2023b, a, Tarzanagh et al., 2023, Valeriani et al., 2023].
@@
Structure in Hidden Representations. Emergence of regularity in last-layer representations during training was noticed in Neural Collapse [Papyan et al., 2020]. Motivated by this phenomenon, subsequent works examined structure of hidden representations and presented theoretical results describing the emergence of Neural Collapse [Wang et al., 2024a, Parker et al., 2023, Zangrando et al., 2024, Garrod and Keating, 2024, Wang et al., 2024b, Hoyt and Owen, 2021, Arous et al., 2023, Zarka et al., 2021, Ben-Shaul and Dekel, 2022, Papyan, 2020, Súkeník et al., 2023]. Our work studies the representations of all layers in LLMs in a local sense through individual trajectories, and uncovers regularity in the representations in LLMs.
@@ 7 Limitations
A main limitation of the scope of our analysis is that due to the significant compute resources required to train LLMs of the size studied in this paper, our exploration is restricted to pretrained LLMs and those fine-tuned on them. Several of these models are trained independently of each other, without carefully controlling for different experimental parameters. Consequently, direct comparisons between models become challenging, and so the cause of varying levels of regularity between LLMs remains unclear. Nonetheless, this makes the consistent emergence of these properties interesting for further study.
@@ 8 Conclusion
A detailed description of the complexities of transformer blocks in the context of language modelling is an active area of research. Our primary goal was to enhance the understanding of the mechanics underlying transformer architectures. We provided a quantitative description of embedding trajectories within LLMs. In a large collection of models, independently trained by various organizations, we discover a remarkable level of regularity among transformer representations characterized by the alignment of residual Jacobians and emergence of linear and exponentially growing trajectories. We observe a correlation between the strength of alignment of residual Jacobian singular vectors and LLM benchmark scores, highlighting the significance of (TA2) and it’s implications regarding performance. These findings open avenues for future research for a deeper understanding the connections between regularity of hidden representations and model specifications, LLM architecture, and generalization.
@@ A.1 Suite of Large Language Models and Prompt Data
We evaluate 38 total LLMs (24 base LLMs and 14 fine-tuned) on 6 datasets from the HuggingFace Open LLM leaderboard [Beeching et al., 2023].
@@
Falcon (40B, 7B) [Almazrouei et al., 2023], Llama-3 (70B, 8B) [AI@Meta, 2024], Llama-2 (70B, 13B, 7B) [Touvron et al., 2023], MPT (30B, 7B) [Team, 2023a, b], Mistral v0.1 [Jiang et al., 2023], Gemma (7B, 2B) [Team et al., 2024], Gemma 1.1 (7B, 2B), NeoX (20B) [Black et al., 2022], Neo (2.7B) [Black et al., 2021, Gao et al., 2020], Pythia (6.9B).[Biderman et al., 2023], and GPT-2 (1.5B, 774M, 355M, 117M) [Radford et al., 2019].
@@
CodeLlama (34B, 13B, 7B) [Rozière et al., 2024], CodeLlama Instruct (34B, 13B, 7B) [Rozière et al., 2024], Mistral-v0.1 Instruct (7.3B) [Jiang et al., 2023], Mistral-v0.2 Instruct [Jiang et al., 2023], CodeGemma [Team et al., 2024].
@@
ARC [Clark et al., 2018], HellaSwag [Zellers et al., 2019], MMLU [Hendrycks et al., 2021], Truthful QA [Lin et al., 2022], WinoGrande [Sakaguchi et al., 2019].
@@
Table 1: LLMs featured in the experiments throughout paper. Included in the table is the parameter budget of each model, the embedding dimension, the number of training tokens, and the Open LLM leaderboard [Beeching et al., 2023] benchmark score.
Model Parameters Layers ($L$) Dim. ($d_{\text{model}}$) Training Tokens Score
Llama-3 70 B 80 8192 15 T
8 B 32 4096 15 T 62.35
Llama-2 70 B 80 8192 2 T 66.05
13 B 40 5120 2 T 55.69
7 B 32 4096 2 T 50.97
Codellama 34 B 48 8192 2 T 55.33
13 B 40 5120 2 T 45.82
7 B 32 4096 2 T 39.81
Codellama (IT) 34 B 48 8192 2 T 43.0
13 B 40 5120 2 T 37.52
7 B 32 4096 2 T 40.05
Orca 13 B 40 5120 2 T 58.64
7 B 32 4096 2 T 54.55
Falcon 40 B 60 8192 1 T 58.07
7 B 32 4544 1.5 T 44.17
Falcon (IT) 40 B 60 8192 1 T 43.26
7 B 32 4544 1.5 T
MPT 30 B 48 7168 1 T 66.98
7 B 32 4096 1 T 56.83
Phi 2 B 32 2560 1.4 T 61.33
1.5 B 24 2048 150 B 47.69
1 B 24 2048 54 B
Mistral-v0.1 7.3 B 32 4096 60.97
Mistral-v0.1 (IT) 7.3 B 32 4096 54.96
Mistral-v0.2 (IT) 7.3 B 32 4096 65.71
Gemma 7 B 28 3072 6 T 64.29
2 B 18 2048 6 T 42.75
Gemma-1.1 7 B 28 3072 6 T 30.0
2 B 18 2048 6 T 60.09
Codegemma 7 B 28 3072 6 T 56.73
2 B 18 2048 6 T 32.19
Neo 20 B 44 6144 41.69
2.7 M 32 2560 0.42 36.20
Pythia 6.9 M 32 4096 0.3 T 39.30
GPT-2 1.5 B 48 1600 34.12
774 M 36 1280 32.07
355 M 24 1024 29.87
117 M 12 768 28.53
A.2 Dynamical Motivation

### v2

@@ 5.4 Correlation with Generalization
For each LLM, we measure the average coupling across depth(which we define to be negative mis-coupling) across prompts in the 6 evaluation datasets (Section 4.2), where for each prompt, $m(A_{ll^{\prime},K}^{n})$ is averaged over layers $l,l^{\prime}\in\{1,\ldots,L\}$. We plot the coupling values against the benchmark scores across several LLMs (Figure 1). Our results reveal a positive correlation between coupling and performance benchmark scores, and is more significant than the relationship between other significant model hyperparameters (Figure LABEL:fig:correlation_hyperparams). This observation suggests a compelling relationship between stronger coupling of singular vectors of Jacobians $J_{l}^{t_{1}t_{2}}$ and improved generalization.
5.5 Regularity in Hidden Trajectories
@@
We identify a considerable degree of linearity in the hidden trajectories of all featured LLMs. The LSS of the trajectories has an average value $4.25$ in trained models, while taking average values of $6.54$ at initialization, and is supported by the low variation across benchmark prompts (Figure 13). Linearity increases with training at varying depths of Pythia 12B (Figure LABEL:figs:lss-training). Linear and expansive behavior of the representations is demonstrated in Llama-3 70B, Mistral 30B, and NeoX 20B through low dimensional projections of embedding trajectories (Figure LABEL:figs:pca).
@@
Figure 13: Linearity of Trajectories. The figure depicts the line-shape score (LSS) of embedding trajectories, as discussed in Section 5.5), computed on 1,200 prompts of the HuggingFace Open LLM Leaderboard (Section 4.2) for a variety of trained (black) and randomly initialized (red) LLMs (Appendix A.1). Plotted are the median values over all prompts, and are accompanied with uncertainty intervals depicting the inter-quartile range of the results for each model. Models are sorted by number of parameters.
5.6 Layer-Wise Exponential Growth
@@
Across all LLMs considered, the hidden trajectories exhibit exponential growth that emerges with training (Figure 14). The majority of the models exhibit a low coefficient of variation across prompts, showing the robustness of this property across a variety of tasks. The expodistance measurements show little dependence on model parameters, and take on consistently low values across all featured LLMs (Figure 14). In contrast, measurements at initialization show equally (rather than exponentially) distanced trajectories, as reflected by the low coefficients of variation in the norms of their layer-wise differences (Figure 22). Under certain assumptions, exponential spacing is justified in Section 6.2.
@@
Figure 14: Coefficient of variation of layer-wise exponent. The figure presents the coefficient of variation of the layer-wise exponent (Section 3.3) computed on 1,200 prompts of the HuggingFace Open LLM Leaderboard (Section 4.2) for trained LLMs (Appendix A.1). Plotted are the median values over all prompts, and are accompanied with uncertainty intervals depicting the inter-quartile range of the results for each model. LLMs are sorted by increasing benchmark performance, while the vertical axis has a base 10 logarithmic scale.
@@ 6.1 Correlation with Generalization
As discussed in Section 5.4, stronger coupling across Jacobian singular vectors is related to improved generalization (Figure 1). This relationship may assist in evaluating the training of transformers. In particular, this finding suggests that the development of training methods to amplify coupling across transformer blocks may lead to favourable model performance.
6.2 Emergence of Exponential Spacing with Coupling
The emergence of exponential spacing in many LLMs can be viewed as a result of coupling under some assumptions on the homogeneity of the singular values across depth. Strong coupling suggests that the linearizations of hidden trajectories satisfy a linear difference equation (see Appendix A.4), given by
$x^{l+1}-x^{l}=Ax^{l}$
Considering the singular value decomposition of $A$ in Equation 18 as $A=U\Sigma V^{T}$, and assuming that there is coupling ($U=V$), the difference equation can be diagonalized in the same basis across depth. Further assuming that the top singular values of the Jacobian blocks display low variation across depth, (Figure LABEL:fig:svalues_trained), we expect
$||x^{l}||\approx e^{\sigma l}\,||x^{0}||,$ (16)
@@ 6.2 Emergence of Exponential Spacing with Coupling
where $\sigma$ is the top singular value. Since we assume $\sigma$ is independent of depth, the above norm is expected to be exponential in $l$, and offers an explanation to the measurements in Figure 14.
@@ 7 Related Work
Residual Networks. ResNets (He et al., 2016) have been viewed as an ensemble of shallow networks (Veit et al., 2016), with studies delving into the scaling behaviour of their trained weights (Cohen et al., 2021). The linearization of residual blocks by their Residual Jacobians was first explored by Rothauge et al. (2019), who examined Residual Jacobians and their spectra in the context of stability analysis, and later by Li & Papyan (2023) who discovered Residual Alignment. Coupling of Jacobian singular vectors in LLM transformers extends previous results for classifier Resnets (Li & Papyan, 2023). We show coupling in $J_{l}^{t_{1}t_{2}}$ across various tokens, which is specific to tokenization in transformers, in addition to demonstrating coupling of $J_{l}^{t_{1}t_{2}}$ across $l$, which was also identified analogously in ResNets (Li & Papyan, 2023). Our measurements quantify the emergence with training of linearity and spacing, including at different depths, which was previously observed mostly qualitatively. Further comparison is included in Appendix A.3.
@@
Neural ODE. Introduced by (Chen et al., 2018), Neural ODE view ResNets as a discretized dynamical process, with past work (Sander et al., 2022) showing the convergence of Residual Networks to a system of linear ODE. (Zhong et al., 2022; Li et al., 2021) studied neural ODE specific perspective to transformers and other LLM architectures. In our work, the presence of residual connections and emergence of Jacobian Alignment in transformers suggests that a discretization of a simple ODE emerges in LLMs.
@@
Information Processing in LLMs. Previously, Katz & Belinkov (2023); Bietti et al. (2023) have described memory and semantic flow in intermediate states of transformers while van Aken et al. (2019) have examined BERT hidden trajectories. Investigation of LLMs for reasoning tasks and interpretability remains an important research focus (Huang & Chang, 2023).
In-context learning. Previous works have observed that LLMs are able to perform in context learning (von Oswald et al., 2023; Bai et al., 2024; Ahn et al., 2023; Akyürek et al., 2023; Xie et al., 2021; Hahn & Goyal, 2023; Xing et al., 2024): the model is capable of performing a particular machine learning task through a few examples provided in a single prompt or context. Their results consider a single layer of trained self-attention as implementing a single step of a gradient-descent-like learning rules across transformers depth, minimizing the MSE of a linear model
$x_{\text{final}}=\min_{x}||Ax-b||^{2}$
A gradient-descent learning rule would cause depth-wise updates to representations of the form
$x_{t+1}=x_{t}+\epsilon A^{T}(Ax_{t}-b)=(I+\epsilon A^{T}A)x_{t}-{\epsilon}A^{T}b$
We should expect similarity in the matrices $I+{\epsilon}A^{T}A$ across depth, which is suggested by coupling.
@@
Hidden Representation Dynamics. Previous studies (Gai & Zhang, 2021; Haber & Ruthotto, 2017; Ee, 2017), have focused on interpreting deep neural networks from a dynamical systems perspective. Throughout training, numerical experiments have shown that both plain networks and ResNets are aptly approximated by a geodesic curve in Wasserstein space (Gai & Zhang, 2021). Other interpretations involve partitioning activation space into basins of attraction, with fixed points representing discrete thoughts (Nam et al., 2023). For further related works, see (Geshkovski et al., 2023b; a; Tarzanagh et al., 2023; Valeriani et al., 2023).
@@
Structure in Hidden Representations. Emergence of regularity in last-layer representations during training was noticed in Neural Collapse (Papyan et al., 2020). Subsequent works examined structure of hidden representations and presented theoretical results describing the emergence of Neural Collapse (Wang et al., 2024a; Parker et al., 2023; Zangrando et al., 2024; Garrod & Keating, 2024; Wang et al., 2024b; Hoyt & Owen, 2021; Arous et al., 2023; Zarka et al., 2021; Ben-Shaul & Dekel, 2022; Papyan, 2020; Súkeník et al., 2023). Within LLMs, recent works have identified uniform structure in token means (Wu & Papyan, 2024), and low dimensional structure across hidden trajectories (Sarfati et al., 2024). Our work studies the representations of all layers in LLMs in a local sense through the interactions of tokens through their Jacobians.
@@ 8 Limitations
A limitation of the scope of our analysis is that due to the significant compute resources required to train LLMs of the size studied in this paper, our exploration is restricted to pretrained LLMs and those fine-tuned on them. Several of these models are trained independently of each other, without carefully controlling for different experimental parameters. Consequently, direct comparisons between models become challenging, and so determining the cause of varying levels of regularity between LLMs becomes difficult. Nonetheless, this makes the consistent emergence of these properties interesting for further study.
@@ 9 Conclusion
Our primary goal was to contribute to the understanding of the mechanics underlying transformer architectures through an analysis of the trajectories of token embeddings and their interactions. Our research builds on the understanding of transformer architectures by revealing the coupling, across depth and token, of singular vectors in the Jacobians of transformer blocks for multiple LLMs trained by various organizations. We additionally observe and quantify the emergence of linearity and exponential spacing with training in the embedding trajectories of LLMs. Notably, we establish a correlation between the strength of this coupling and benchmark performance on the HuggingFace Open LLM Leaderboard, highlighting the significance of transformer block coupling for generalization. These findings open avenues for future research, encouraging deeper exploration into the connections between regularity of hidden representations, model specifications, and generalization.
@@ A.1 Suite of Large Language Models and Prompt Data
We evaluate 38 total LLMs (24 base LLMs and 14 fine-tuned) on 6 datasets from the HuggingFace Open LLM leaderboard (Beeching et al., 2023).
@@
Falcon (40B, 7B) (Almazrouei et al., 2023), Llama-3 (70B, 8B) (AI@Meta, 2024), Llama-2 (70B, 13B, 7B) (Touvron et al., 2023), MPT (30B, 7B) (Team, 2023a; b), Mistral v0.1 (Jiang et al., 2023), Gemma (7B, 2B) (Team et al., 2024), Gemma 1.1 (7B, 2B), NeoX (20B) (Black et al., 2022), Neo (2.7B) (Black et al., 2021; Gao et al., 2020), Pythia (6.9B).(Biderman et al., 2023), and GPT-2 (1.5B, 774M, 355M, 117M) (Radford et al., 2019).
@@
CodeLlama (34B, 13B, 7B) (Rozière et al., 2024), CodeLlama Instruct (34B, 13B, 7B) (Rozière et al., 2024), Mistral-v0.1 Instruct (7.3B) (Jiang et al., 2023), Mistral-v0.2 Instruct (Jiang et al., 2023), CodeGemma (Team et al., 2024).
@@
ARC (Clark et al., 2018), HellaSwag (Zellers et al., 2019), MMLU (Hendrycks et al., 2021), Truthful QA (Lin et al., 2022), WinoGrande (Sakaguchi et al., 2019).
@@
Table 1: LLMs featured in the experiments throughout paper. Included in the table is the parameter budget of each model, the embedding dimension, the number of training tokens, and the Open LLM leaderboard (Beeching et al., 2023) benchmark score.
Model Param. Layers ($L$) Dim. ($d_{\text{model}}$) Tokens Score Pos.Encoding
Llama-3 70 B 80 8192 15 T RoPE
8 B 32 4096 15 T 62.35 RoPE
Llama-2 70 B 80 8192 2 T 66.05 RoPE
13 B 40 5120 2 T 55.69 RoPE
7 B 32 4096 2 T 50.97 RoPE
Codellama 34 B 48 8192 2 T 55.33 RoPE
13 B 40 5120 2 T 45.82 RoPE
7 B 32 4096 2 T 39.81 RoPE
Codellama (IT) 34 B 48 8192 2 T 43.0 RoPE
13 B 40 5120 2 T 37.52 RoPE
7 B 32 4096 2 T 40.05 RoPE
Orca 13 B 40 5120 2 T 58.64 RoPE
7 B 32 4096 2 T 54.55 RoPE
Falcon 40 B 60 8192 1 T 58.07 RoPE
7 B 32 4544 1.5 T 44.17 RoPE
Falcon (IT) 40 B 60 8192 1 T 43.26 RoPE
7 B 32 4544 1.5 T RoPE
MPT 30 B 48 7168 1 T 66.98 ALiBi
7 B 32 4096 1 T 56.83 ALiBi
Phi 2 B 32 2560 1.4 T 61.33 RoPE
1.5 B 24 2048 150 B 47.69 RoPE
1 B 24 2048 54 B RoPE
Mistral-v0.1 7.3 B 32 4096 60.97 ALiBi
Mistral-v0.1 (IT) 7.3 B 32 4096 54.96 ALiBi
Mistral-v0.2 (IT) 7.3 B 32 4096 65.71 ALiBi
Gemma 7 B 28 3072 6 T 64.29 RoPE
2 B 18 2048 6 T 42.75 RoPE
Gemma-1.1 7 B 28 3072 6 T 30.0 RoPE
2 B 18 2048 6 T 60.09 RoPE
Codegemma 7 B 28 3072 6 T 56.73 RoPE
2 B 18 2048 6 T 32.19 RoPE
Neo 20 B 44 6144 41.69 RoPE
2.7 B 32 2560 0.42 36.20 Sine
Pythia 12 B 36 5120 0.3 T 58.9 RoPE
6.9 B 32 4096 0.3 T 39.30 RoPE
2.8 B 32 2560 0.3 T RoPE
1.4 B 24 2048 0.3 T RoPE
1 B 16 2048 0.3 T RoPE
GPT-2 1.5 B 48 1600 34.12 Sine
774 M 36 1280 32.07 Sine
355 M 24 1024 29.87 Sine
117 M 12 768 28.53 Sine
A.2 Additional Metrics
A.2.1 Visualization of Trajectories with PCA
A.3 Comparison to ResNets
Coupling. Our results complement and build upon those of (Li & Papyan, 2023), who have observed the coupling of singular vectors of Residual Jacobians in classification ResNets. We observe coupling across depth in a wide range of LLMs, where Jacobians are evaluated at the sequence of embeddings, with respect to the current token, whereas in RA Jacobians are evaluated at a single representation. Additionally, with transformers we may analyze coupling not only across depth but also across tokens. We observe coupling in LLMs across tokens in a variety of ways. Further, we consider and analyze the relationship between coupling and generalization.
Linearity and Equidistance. Linearity in hidden trajectories, as observed in ResNets Li & Papyan (2023); Gai & Zhang (2021), also emerges with training in LLMs. The mean LSS value among the evaluated LLMs is $4.24$ (Figure 13), greater than LSS measurements observed for ResNets (Gai & Zhang (2021), page 18) which range between $2.0$-$3.0$ (due to varying trajectory length and hidden dimension). In both architectures, training induces improved linearity and regularity (Figure LABEL:figs:pca) in trajectories. In contrast to ResNets, trajectories are not equidistant, instead showing exponential growth between layers (Figure 14). We quantify this spacing through a low coefficient of variation, displaying the presence of exponential growth in token trajectories. In both classifier ResNets and LLM transformers, there is an evident level of regularity in hidden representations (Figure LABEL:figs:pca).
Rank of Jacobians. Li & Papyan (2023) show that residual Jacobians have rank at most $C$, the number of classes. This analogous result automatically holds for LLMs since the vocabulary size is significantly greater than the embedding dimension of the transformer blocks.
Singular Value Scaling. Li & Papyan (2023) observe that top singular values of residual Jacobians scale inversely with depth. In trained LLMs, however, top singular values do not show a consistent depth scaling across models (Figure LABEL:fig:svalues_trained), notably differing from classification ResNets. In addition, the distribution of singular values at each layer varies significantly between models. Singular value scaling is more present in untrained transformers (FigureLABEL:fig:svalues_untrained), likely caused by additional layer normalizations in residual blocks (Section 4.1).
A.4 Dynamical Motivation